OpenStudy (anonymous):

Find the equation for the normal line to f at the point (-2, f(-2)). f(x)= -x +4x^2

OpenStudy (tkhunny):

Have you considered the 1st Derivative as a clue to find the slope of the Normal Line?

OpenStudy (anonymous):

the first derivate is 8x

OpenStudy (tkhunny):

Evaluated at x = -2. What is the slope of the tangent line at that point?

OpenStudy (anonymous):

-16?

OpenStudy (tkhunny):

Good. Now, what is the slope of the NORMAL line at that point?

OpenStudy (anonymous):

18!

OpenStudy (anonymous):

sorry 14?

OpenStudy (anonymous):

so is y -18 = 16x?

OpenStudy (tkhunny):

?? The Normal is Perpendicular to teh Tangent. Give it another go.

OpenStudy (anonymous):

wait!, okay the derivative for f(x) = -x +4x^2 is equal to 8x-1, thus the answer is -17. okay after it whats next?

OpenStudy (anonymous):

0.o?

OpenStudy (anonymous):

Your question is not making sense? Or am i reading it wrong.

OpenStudy (tkhunny):

Whoops! Missed the x. Good catch. f(x) = -x + 4x^2 f(-2) = 2 + 4(4) = 2+16 = 18 f'(x) = -1 + 8x f'(-2) = -1 - 16 = -17 The tangent line at (-2,18) is (y-18) = -17(x+2) The normal line is perpendicular to the tangent line and passes through the same point. You tell me what that equation is.

OpenStudy (anonymous):

answer*** not question.

OpenStudy (anonymous):

where the (x+2) come from?

OpenStudy (tkhunny):

You should know the Point-Slope form of a line. Point (a,b) Slope m (y-b)=m(x-a) Don't forget your algebra!!

OpenStudy (anonymous):

M.A.T.H stand for Mental Abuse to Humans, my brain does not have enough memory for more!

OpenStudy (tkhunny):

Blah, blah. You can do it. Hang in there! :-)